### v1

# A search for Galactic post-AGB stars in Gaia DR3††thanks: Tables A6 and A7 are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A

Abstract
Context. When low and intermediate-mass stars leave the Asymptotic Giant Branch (AGB) phase, and before they reach the Planetary Nebulae (PN) stage, they enter a very brief and rather puzzling stellar evolutionary stage named post-AGB. The post-AGB phase has a very short duration, of the order of a few thousand years at most, so the number of objects confirmed in this phase is really small, and thus, our understanding of this elusive stellar evolutionary stage is very constrained.
Aims. To provide a reliable catalogue of galactic post-AGB stars together with their physical and evolutionary properties obtained through Gaia DR3 astrometry and photometry. As an added product, we provide information for a sample of other types of stellar objects, whose observational properties mimic those of post-AGB stars.
Methods. Post-AGB stars are characterised by their infrared excesses and high luminosities. The publication of precise parallaxes in Gaia DR3 made it possible to calculate accurate distances and to revise the derivation of luminosities for post-AGB candidates to discard objects outside the expected luminosity range. We started by identifying post-AGB stars or possible candidates from the bibliography and, then, searched for their Gaia DR3 counterpart sources. Using the available photometry, interstellar extinction, literature spectroscopically derived temperatures or spectral types and parallax-derived distances, we fitted their Spectral Energy Distributions (SED) and we estimated their luminosities and circumstellar extinctions. When compared to models, luminosity values allowed us to disclose objects that are likely post-AGB stars from other target types. Their position on the Hertzsprung-Russell (HR) diagram allows direct comparison with updated post-AGB evolutionary tracks and an estimation of their masses and evolutionary ages.
Results. We obtained a sample of 69 reliable post-AGB candidates that meet our classification criteria, providing their coordinates, distances, effective temperature, interstellar and circumstellar extinction, luminosity, mass, and evolutionary age. In addition, similar data for other stellar objects in our initial compilation, such as supergiant stars and Young Stellar Objects (YSOs), is provided. Our identifications and parameters are compared with others found in recent literature about the subject.
Conclusions. We have filtered out the data that have the best precision in parallaxes and distances to obtain more accurate luminosities, which consequently, allows us to classify with confidence the objects of the sample among different stellar phases. In turn, this allows us to provide a small but reliable sample of post-AGB objects. Derived mean evolutionary time and average mass values are in agreement with theoretical expectations, and with the mass mean value obtained in a previous work for the subsequent evolutionary stage, the Planetary Nebula stage.
Key Words.:
Astronomical data bases – Stars: AGB and post-AGB, distances, fundamental parameters, Hertzsprung-Russell and C-M diagrams – Virtual Observatory tools1 Introduction
The stellar phase known as post Asymptotic Giant Branch (post-AGB) is a very fast (a few thousand years) and quite unknown phase, that takes place at the end of the low and intermediate-mass star lifetime, after the AGB phase and before the Planetary Nebula (PN) stage, where the star ionises the previously ejected envelope. There is no exact definition for the beginning of the post-AGB phase, and its onset depends on the stellar mass and metal content. Furthermore, the departure from the AGB phase is defined somewhat arbitrarily in the stellar evolution models. For Vassiliadis & Wood (1993) the AGB mass-loss is terminated when the envelope mass is reduced to such a value that the stellar effective temperature increases beyond a reference value by an amount of log(Teff)=0.3. Other authors have different criteria. In Miller Bertolami (2016), they set the onset when, due to the stellar winds, the H-rich envelope mass drops below 1% of the stellar mass.
During the post-AGB phase, the star will evolve at almost constant luminosity towards hotter effective temperatures, while its envelope will expand into the interstellar medium. Due to the high stellar temperature, this envelope composed of gas and dust will start to be ionised (see e.g. Villaver et al. 2002). At this moment, the star will enter the proto-planetary nebulae phase. As in González-Santamaría et al. (2021) we can assume a minimum stellar temperature of 13 000 K for a transition stage from a preplanetary nebula (Weidmann et al. 2020) to about 24 000 K for a complete ionisation of the nebula (Kwok, 2000).
Over the last years, several studies have been carried out to identify post-AGB stars to better characterise this brief stellar phase. Garcia-Lario et al. (1997), based on their IRAS infrared colours, identified 110 possible post-AGBs, most of them without optical counterparts. At the beginning of this century, Suárez et al. (2006) provided optical counterparts for more than 100 post-AGB candidates, previously proposed as such based on their IRAS fluxes. Almost in parallel, a more extensive catalogue of post-AGB candidates was presented by Szczerba et al. (2007), which was later extended in Szczerba et al. (2012). This is known as the Torun catalogue of post-AGB stars, which is the largest one to date containing 296 sources classified as either Likely or Possible post-AGBs. The sample of post-AGB stars has been expanded to the Large Magellanic Cloud (LMC) by Kamath et al. (2015), including also post-RGB (Red Giant Branch) stars. These stars are believed to be produced from binary interactions in the RGB phase, and their observational properties mimic those of post-AGB stars, although they are not expected to reach luminosities as high as those of post-AGB stars (Kamath et al. 2015). In a previous article (González-Santamaría et al. 2021), we found that a fraction close to 50% of the central stars of PNe are red, and therefore possibly unresolved binary systems. Finding the relative number of stars of one type and another, single or binary, at those evolutionary stages, can give us clues about the fraction of binary stars with evolved companions.
Recent work by Kamath et al. (2022) studies the evolutionary state of 31 galactic post-AGB candidates with chemical abundance information. By using Gaia EDR3 distances and known photometry, they build the Spectral Energy Distribution (SED) for these objects, and by fitting them to models, they estimate their luminosities and temperatures. It is important to note that 20 of the objects analysed have quite poor astrometry in Gaia DR3.
In Parthasarathy et al. (2020), the properties of 8 post-AGB candidates were analysed making use of Gaia DR2 astrometric and photometric data. Also using Gaia, in this case DR3, Oudmaijer et al. (2022) investigate the nature as possible post-AGB of 249 objects. It is noteworthy that most of the objects here selected present important uncertainties in their parallaxes, which leads to very unreliable distance values. This work considers good parallaxes those with relative error () between 10% and 100%. Additionally, we detected some inconsistencies in this study that we shall comment on in section 3.
Finally, in a recent work, Aoki et al. (2022) studied the evolutionary state of 20 post-AGB candidates by using Gaia DR2 and EDR3.
We aim to go further into the identification and analysis of the properties of bona fide post-AGB stars by making a selection of objects with accurate astrometric measurements in Gaia DR3, thus allowing a precise location in the Hertzsprung-Russell (HR) diagram and, consequently, a quite reliable classification as post-AGB objects. Selecting among Galactic post-AGB stars candidates with good astrometry in Gaia DR3, in a first approximation, would exclude most astrometric binaries from it. The general rule is that the threshold of Gaia astrometric quality parameter Renormalised Unit Weight Error (RUWE) 1.4 is used to indicate single well-behaved solutions (Lindegren et al. 2018, 2021). Sources whose observations are inconsistent with Gaia astrometric 5-parameter model could be caused by binarity (Lindegren et al. 2018) or other factors that cause the photocentre of the source to wobble during the Gaia observation window. In summary, a restrictive criterion in the astrometric quality will help us to select objects with a higher probability of being individual sources, and on the other hand, it will help us to better estimate their luminosity and evolutionary stage. The incidence of binaries in our resulting sample, however, will be addressed by studying the SED morphology and also in the literature context.
We started by collecting a sample of post-AGB candidates as complete as possible from the currently available catalogues (Sect. 2) and we implemented restrictive filtering over the astrometric quality of the objects, to only keep those with the most accurate distance values. Concretely, we gathered a general sample of 964 post-AGB candidates from the literature of which we filtered out a subset of 178 objects with accurate astrometric measurements in Gaia DR3. A good distance determination is not enough to obtain a reliable adjustment of the SED; it is necessary to have additional information about the temperature and/or extinction in the direction of the source, allowing to consistently derive the luminosity of the object. To better constrain the value of the total extinction, interstellar and circumstellar, in the direction of the source, we opted to limit our work to those objects with available interstellar extinction measures in the literature. From the previous sample of 178 objects, we kept 146 galactic post-AGB candidate stars with literature values of their interstellar extinction. In this last sample, the information available in the Simbad database as well as images of every sky field in Aladin Sky Atlas (Bonnarel et al. 2000) were analysed. 28 sources were finally discarded for further analysis because either they have been already classified as PNe (5 objects), or they have effective temperatures in the literature higher than 24 000 K (11 objects), or the identification of the optical counterpart is dubious (12 objects). Our final working set turned out to be 118 objects. We found that for about 67% of them, temperature values were derived from spectral analysis, with a fraction of them coming from high-resolution spectral analysis. For the rest, average temperatures corresponding to their spectral types were used. Section 3.2 describes in detail the problems associated with the different quality of temperature determinations used in our work.
Photometry in a wide spectral region compiled by the spanish Virtual Observatory SED Analyser (VOSA111http://svo2.cab.inta-csic.es/theory/vosa) alowed us to build the SED for each object. The knowledge of distances and temperatures allowed us to obtain luminosities and to estimate the total extinction from the SED fitting (Sect. 3.3). We have used current knowledge of interstellar extinction in the direction of every object to verify the consistency of the total extinction obtained in the fits.
Based on evolutionary tracks, we used the distribution of objects in an HR diagram to confirm 69 objects out of 118 candidates as post-AGB stars. Meanwhile, some other objects could be classified as Horizontal Branch stars (3 objects), luminous supergiants (3 objects), Young Stellar Objects (YSO –14 candidates), and the rest would remain as unconfirmed post-AGB candidates, –details to be given in Sec. 3.5. In Section 3.6, this classification is compared with those presented in other recent papers about the subject. In Sect. 4, we focussed on analysing the evolutionary properties of the set of 69 objects that we identify as belonging to the post-AGB stage. In Sect. 5, we comment on some interesting objects, and, finally, in Sect. 6, we summarise our conclusions.
2 Sample selection: methodology
The first step in this research was to collect all the objects from the literature catalogued as confirmed or possible, post-AGB stars. For this purpose, we used the online222https://fox.ncac.torun.pl/camkweb/postagb2.php Torun catalogue of Galactic post-AGB stars (Szczerba et al. 2012), the Simbad astronomical database and the spectroscopic atlas of post-AGB and planetary nebulae by Suárez et al. (2006) (from now on Suarez et al. catalogue).
Concretely, we gathered all objects catalogued as Likely (209) or Possible (87) in the Torun catalogue, as Post-AGB (331) or post-AGB Candidate (507) in the Simbad database and as Post-AGB (102) in the Suarez et al. catalogue. To obtain the union of the three catalogues, firstly, we matched the Torun objects with those in Simbad using a 1 arcsec cross-match radius. As a result, we obtained a set of 929 objects. Subsequently, we matched those with Suarez et al. catalogue (by the same method and using coordinates from Torun when possible), obtaining a final sample of 964 objects.
The next step was to cross-match our list of objects with the Gaia DR3 archive to obtain their parallaxes, and consequently their distances. We used again a searching radius of 1 arcsec from the literature coordinates. Note that in some cases, the coordinates from Simbad or Suárez et al. catalogues differ slightly from the coordinates from the Torun catalogue; this can lead to discrepancies in the identification of the Gaia DR3 source. We decided to prioritise the Torun coordinates, as they come from a revised compilation of post-AGB stars. As a result, we were able to identify 843 objects as Gaia DR3 sources, around 87% of the whole sample. Finally, we discarded those objects with unknown distances in Bailer-Jones et al. (2021), and we ended up with 765 objects.
In Gaia DR3, parallaxes () show a bias or zero point () that should be considered and subtracted from the measured value to obtain the actual parallax (). According to Lindegren et al. (2021) this zero point has a mean value of -17 as, although its value varies depending on the star celestial position, colour and magnitude. Gaia web provides a Python code to estimate this zero point333https://gitlab.com/icc-ub/public/Gaiadr3_zeropoint. After obtaining the zero points, we corrected the parallaxes with the following simple expression:
Uncertainties in Gaia parallaxes include both the value given in Gaia DR3 archive, the internal uncertainty, as well as a systematic uncertainty that depends on the source brightness. We followed the prescriptions given in Fabricius et al. (2021) to estimate the parallax total uncertainties of our objects.
From the inverse of the parallaxes, it is possible to estimate the stellar distances, but this simple approach is only valid for objects with relative uncertainties as low as 10%. In general, it is not our case, so we used the distances calculated by a Bayesian approach for the Milky Way stars by Bailer-Jones et al. (2021), which consists of assuming an a priori probability volume density of stars in the Galaxy that decreases exponentially on an appropriate distance scale. This method not only provides an estimated distance for a source but also gives low and high error distance bounds.
To infer useful stellar properties that depend on distance, such as luminosity, it is important to have precise distances to these objects. So we decided to apply filtering to our sample according to the distance uncertainties as well as to the astrometric quality. Astrometric quality indices such as UWE (Unit Weight Error, see Gaia official website) and RUWE (Renormalised Unit Weight Error) beyond certain boundaries prevent Gaia users from errors in the astrometric solution that can be due to binarity or to irregularities in the fitted source. We decided to use the same filtering criteria as that used in one of our previous works, where we did a similar study but over the galactic sample of planetary nebulae central stars (González-Santamaría et al. 2021). This filtering consists of the following constraints: relative error in parallax and distance (for both lower and upper bounds) below 30% and astrometric quality parameters UWE and RUWE below certain threshold values recommended in Gaia documentation (UWE¡1.96 or RUWE¡1.4). After applying these constraints, we ended up with a set of 178 objects that we consider to have good astrometric measurements in Gaia DR3.
If we analyse the galactic distribution of these stars we appreciate a small cluster of 21 objects in a region of mid-south latitudes (see Fig. 1) at longitudes coincident with those of the direction of the LMC (Large Magellanic Cloud). Furthermore, we verified that all of these objects are included in the van Aarle et al. (2011) catalogue of LMC post-AGB stars, while 10 of them are also included in Kamath et al. (2015). So, we decided to exclude those stars from our present study and focus on the remaining 157 Galactic post-AGB candidates.
In Table LABEL:table:candidates we provide general data of the objects in this later sample: a running number, name, Gaia DR3 ID, J2000 coordinates, G magnitude, interstellar extinction in the V-band, spectral type, reference for identification as post-AGB (Torun, Simbad and Suárez et al. catalogues), and a flag providing information about binarity, variability and the reason for excluding the source for further analysis, as it was mentioned in the previous section, together with a reference.
The distance distribution of all these post-AGB candidates is shown in Fig. 2. We will not attempt to do any analysis of the completeness of the sample since it presents several observational biases, among which is the impossibility of detecting the optical counterpart of some of the infrared sources with data in the IRAS catalogue.
3 Galactic post-AGB stars identification
The luminosity of a star is a very useful property to distinguish between post-AGB stars and other stellar objects with similar colours, that are located out of the Main Sequence (MS) of the HR diagram, as it occurs with YSOs. To narrow the luminosity range that corresponds to post-AGB stars, we resorted to different evolutionary models for hydrogen-burning post-AGB stars. Concretely, those classical ones by Vassiliadis & Wood (1993) (for masses between 1 and 5 ), and by Bloecker (1995) (for masses between 1 and 7 ), and more recently by Miller Bertolami (2016), which includes different metallicities (for masses between 0.8 and 4 ). Although the range of masses of the progenitor stars is different in each of the models, it should be noted that the evolution of stars with masses greater than 4 is very fast, so rather poor statistics will be expected for objects of these masses and higher. Their expected number should also be very small given the initial mass function. We found a reasonable agreement (see Figure 3) among the models that the range of luminosities that correspond to this evolutionary phase should be .
Such luminosity range for post-AGB stars meets the criteria used by Kamath et al. (2015), who consider that the luminosity range for post-AGB stars is between 2500 and 35,000 . According to these authors, objects above such upper limit may be supergiants or hypergiants, high-mass stars that quickly initiate helium core fusion after they have exhausted their hydrogen, and continue fusing heavier elements after helium exhaustion until they develop an iron core, at which point the core collapses to produce a Type II supernova. In contrast, objects below the lower limit may be post-RGB stars, YSOs, or eventually, other evolved stars such as Horizontal Branch (HB) stars.
The main problem with classifying high luminosity stars as either post-AGB evolved objects or high luminosity massive objects is that both types of objects share many observational features: similar optical spectra, unstable and extended atmospheres, expanding gas-dust envelopes, high IR excesses, and similar IRAS colours (Garcia-Lario et al. 1997). This matter is dealt with in detail by Klochkova & Chentsov (2018), who argue the need to determine and compare various parameters: position in the Galaxy, luminosity, wind parameters, SED, and chemical composition to allow for accurate classification. For this reason, the few objects that we found up the limit will be discussed individually (see Sect. 5).
A different problem occurs with objects below the minimum luminosity indicated by post-AGB evolution models, those with . As we mentioned before, in Kamath et al. (2015) post-AGB candidates in the Magellanic Clouds located in such range were tentatively identified as post-RGB stars. According to these authors, these stars are most likely the result of binary interaction in which their evolution towards the AGB is interrupted, but only in some cases their binary nature has been confirmed, and it could also be the case that some of them are the result of a merging process. It is beyond the scope of this paper to analyse such objects and, for this reason, we would refer to them as unconfirmed post-AGB candidates.
Thus, to classify our objects, as post-AGB stars or another type of stellar objects, it is necessary to calculate accurate luminosities and also to have a reliable estimation of their temperatures.
3.1 Interstellar reddening
To explore the best possible determination of the interstellar extinction values for our objects, we used Gaia DR3 coordinates and distances and searched the bibliography for the corresponding extinction values. After analysing and contrasting data from different catalogues and dust maps, we decided to use the extinction values from Stassun (2019). Concretely, we obtained values from this catalogue by performing a cross-match between our objects sample and the ’TESS Input Catalog v8.0’ (Stassun 2019), using the Topcat tool (Taylor 2005). The Stassun catalogue contains extinctions from dust maps for 146 objects of our sample. We always used distance-dependent extinction values when they were available. Extinction errors are provided for 116 objects, and around 80% of them are below magnitudes. Interstellar extinction values are listed in Table LABEL:table:candidates.
3.2 Effective temperatures
We searched the Simbad database for temperature values and references for each object in our sample of 146 candidates and, as a result, we discarded 28 objects that were too hot or with a dubious identification of the central source in the literature. In consequence, we ended up with a sample of 118 objects, that we will consider our final sample. Details are included in Table LABEL:table:candidates. Regarding temperature values, very different situations were found. For approximately 67% of objects, there are precise temperature determinations, which come from spectral analysis. This is the case of 11 objects in common with the work of Kamath et al. (2022), or those in common with Corporaal et al. (2023) or Mello et al. (2012) (see references in Tables A.2, A.3, A.4 and A.5).
For several cases, Teffs come from spectral types derived from medium-resolution spectra, as in the Suárez et al. (2006) catalogue. For others, only Simbad spectral types are available, some of them covering a quite wide range of subtypes or even types. We have also noticed that for some objects, the spectral classifications in the MK system, generally old, are quite discrepant with the spectroscopic temperatures obtained in more recent publications. For instance, the star BD+48 1220 is assigned a spectral type A4Ia (8550 K) in Simbad, from reference Hardorp et al. (1965), while Ting et al. (2019) reports a value of 6389 K from APOGEE spectra analysis. This leads us to deduce that, at least for some cases, the effective temperatures obtained from spectral types may have inaccurate values.
Following the precision of literature values, temperatures from spectral analysis have been prioritised over temperatures obtained by spectral classification in MK types, which in turn have been prioritised over average temperatures obtained directly from the spectral type in the Simbad database. Tables A.2, A.3, A.4, and A.5 list the effective temperature adopted for each object together with a reference and a flag indicating its origin.
3.3 Luminosity and total extinction from SED fitting
To estimate the luminosity of a star, its bolometric flux or magnitude, distance, and interstellar extinction values are needed. A simple approach consists of obtaining the stellar photospheric magnitude V and then applying the bolometric correction to derive the bolometric magnitude. Alternatively, stellar photometry in several bands can be used to build the SED, and then, by fitting it to a certain model, the stellar temperature and luminosity can be predicted. This simple approach is not possible in most cases because of the presence of dust in the interstellar medium, which reddens the spectral distribution and converts the determination of parameters by fitting with a model into a degenerate problem between temperature and extinction.
We will assume as valid the temperature values obtained from the literature with the methodology explained in the previous section. We then used a procedure that allowed us to obtain the luminosity by fitting the SED, introducing the total extinction necessary to obtain the already known temperature value within a range of 250 degrees, which can be considered an acceptable value for the errors of the temperatures assigned from the literature to each object.
The VOSA software is the spanish Virtual Observatory tool designed, among other uses, to estimate effective temperature (Teff), gravity and luminosity based on stellar photometry. The user provides the coordinates of the source, the source distance and its uncertainties and the system looks for observed flux/magnitudes (and their errors) by querying several photometric catalogues accessible through VO services to achieve a wavelength coverage of the data to be analysed as wide as possible. It is possible then to choose among different stellar models to perform the fitting. Concretely, we choose Kurucz models (Castelli & Kurucz, 2003) because they are well fitted for our range of temperatures and the evolutionary stage of post-AGB stars. The VOSA software then performs the absolute flux calibration of the observational data, using the information regarding the available filters (zero points, transmission curves, etc.).
Next, it determines the synthetic photometry for the models with physical parameters in the range selected by the user (in our case, log[g] values between 0 and 5, and metallicity between -4 and 0.5). Dust extinction is also an input to the system, it is provided with no uncertainty together with the selection of an appropriate extinction law. Concretely, VOSA make use of the extinction law by Fitzpatrick (1999) improved by Indebetouw et al. (2005) in the infrared. Next, the best fitting model is provided by VOSA, together with the derivation of the corresponding stellar parameters: Teff and luminosity (as well as a value for log[g], metallicity and overabundances of -elements with respect to iron).
In the problem that interests us, we carried out an iterative procedure which consisted of providing an input test value of the extinction (starting with a value close to the interstellar extinction value from 3D maps) and checking which temperature value was obtained in the SED fitting. We then modified the extinction value in steps of 0.05 magnitudes until a temperature value closer to the literature one within the uncertainty of 250 K mentioned before was obtained. Note that total extinction values obtained from the SED fit can be compared with those from the extinction maps to derive the contribution of the circumstellar component.
Post-AGB candidates in the Torun catalogue were identified as objects with infrared excesses due to the presence of a dust envelope or a disc. This means that those infrared excesses should be accounted for when fitting the SEDs. In VOSA the excesses are detected by iteratively calculating (adding a new data point from the SED at a time) in the mid-infrared (wavelengths redder than 2.5 ) the parameter as defined in Lada et al. (2006). The theoretical spectral models used by VOSA are based on stellar atmospheres and as a result for the calculation of the errors of the fit, the tool only considers those data points of the SED that correspond to bluer wavelengths than the one where the excess has been flagged.
The VOSA service can be used to build SEDs by querying a large variety of photometric surveys available on the platform. In general, we found fluxes for our objects from several catalogues, covering ultraviolet, visible, and near-infrared wavelengths. The most common flux sources used are: 2MASS, DENIS, IRAS, AKARI, WISE, Tycho, Paunzen, UBV, Gaia DR3, Gaia XPy, Pan-Starrs, and GALEX. The flux ranges are shown in Appendix B444https://zenodo.org/uploads/11569760 figures and the individual fluxes for each object are listed in Table A.6 and Table A.7, only available at CDS. The SEDs we provide have been carefully analysed to check for bad data points and then as mentioned before, fitted to Kurucz models (Castelli & Kurucz 2003). We found these models very suited for our sample, as they cover a large temperature range, from 3500 K to 50,000 K.
This procedure has allowed us to derive coherent temperature/total extinction pairs and the luminosity for each of the 118 stars in our final working sample. The uncertainty values in luminosity were estimated by VOSA tool using the uncertainties in the photometry and taking into account the distance uncertainties (lower and upper limits) that we provided as input. The VOSA software provides uncertainties in luminosities, in general, below 10%, which can be considered as a lower limit. As mentioned, VOSA does not support the use of errors in the extinction values to compute the SED fitting. More information about the fitting procedure is provided in Bayo et al. (2008) and in VOSA documentation 555http://svo2.cab.inta-csic.es/theory/vosa.
This procedure, together with accurate distances from Gaia parallaxes, allowed us to obtain values for the luminosities and the total extinctions in agreement, for instance, with those given by Kamath et al. (2022) as it is illustrated in Table 1. However, the fitting procedure used to estimate the luminosity and temperature values has its own limitations, as it does not take into account the uncertainty of the total extinction values (affecting luminosity uncertainties), and it also depends on the fitting models.
The VOSA tool fitting of Kurucz models allows also to obtain tentative values of log[g] (which covers a range between 0 and 5) and metallicity (which covers a range between -4 and 0.5) for our sample stars, although more reliable values for such parameters can be obtained from spectroscopy when available. In Appendix B we provide the fitted SEDs for all the 118 objects in our final sample, including such parameters. Luminosity values together with lower and upper uncertainties, with the limitations explained before, are presented in Tables 3, 4, 5 and 6.
Following Kamath et al. (2015), we analysed the shape of the SEDs and provided a classification in three different types (stellar, shell, or disc). This can give us some additional clues about the possible incidence of binarity in our sample. According to these authors, disc-type SEDs are related to binarity. We classified 30 SEDs as disk-type. The SED morphological classification for our objects is shown in Tables 3, 4, 5 and 6. We opted to locate the disc-type SEDs in an HR diagram along with the rest of the objects as their identification as binaries is tentative and not, in general, confirmed.
Figure 4 depicts the temperature distribution for our final sample objects. Most stars (84%) have values below 10 000 K, as it is expected for post-AGB stars, while, only three stars exhibit effective temperatures above 20 000 K. These last three are sources with infrared flux excesses which have already started to ionise their envelopes, on their way to the PN phase.
If we compare our temperature determinations with those obtained by Oudmaijer et al. (2022) for the 59 objects in common in both samples, we find (Fig. 5) that the temperatures in Oudmaijer et al. (2022) tend to be slightly higher than the ones we obtained. The mean difference being
The temperatures in Oudmaijer et al. (2022) are based on spectral types collected from the Simbad database, which have very different origins and qualities, which might explain the discrepancies.
Regarding the stellar luminosity, Fig. 6 shows that most of the candidate objects have luminosities between , a region that includes the main luminosity range expected for post-AGB stars, but the histogram includes a wide zone of underluminous objects as well.
Figure 7 shows the comparison between our luminosity values and those presented in Oudmaijer et al. (2022) for the 69 stars in common. Note that in Oudmaijer et al. (2022), not all objects with a luminosity value have a temperature value. The luminosity values from Oudmaijer et al. (2022) are very similar to those obtained in this work, the mean difference being
Oudmaijer et al. (2022) determined their luminosities through the dereddened integrated fluxes obtained from Vickers et al. (2015) and multiplying by the square of the distances from Gaia DR3 parallaxes. The authors indicate that the errors in the fluxes are of the order of 20%, which could explain some differences. It is also important to note that Vickers et al. obtained their integrated fluxes assuming default values for the luminosity which implies then that the values in Oudmaijer et al. (2022) are calculated using a rather circular argument.
3.4 Galactic heights and membership to the Halo
Gaia DR3 precise distances allowed us to calculate the galactic distribution of our final sample. Figure 8 depicts the galactic height as a function of the galactic longitude. It also shows the commonly adopted limits for the main structures in the Milky Way: thin disk, thick disk, and halo.
This distribution allows us to tentatively assign the 118 objects either to the disk 83 objects with kpc) or to the halo (35). Suspected halo stars are flagged with ’H’ in Tables A2, A3, and A5. Although this classification is adopted to compare their position in the HR diagram with evolutionary tracks suited for each of the populations, we are well aware that it will benefit from a spectroscopic confirmation.
3.5 Classification: identification of post-AGB stars and other objects
Once luminosity values and their uncertainties are known, we can apply the luminosity thresholds discussed before. We obtain that 69 objects can be classified as post-AGB stars bona fide candidates, 46 objects can not be confirmed as post-AGB stars because we derive a luminosity lower than 2500 for them, and 3 objects above the high 35,000 luminosity threshold are classified as supergiant stars.
From the sample of 46 stars with luminosities lower than 2500 , 5 are found to be YSOs in Molecular Clouds (see further in this section), 3 are suspected or confirmed to be Horizontal Branch (HB) stars, and 38 remain as unclassified. From this last group, 9 objects with luminosity lower than 100 are tentatively classified as possible YSOs as will be discussed later. The properties of the objects in the main categories are listed in Table A.2 (post-AGB stars), Table A.3 (unconfirmed post-AGB candidates), Table A.4 (YSOs candidates), and Table A.5 (supergiants and HB stars).
To illustrate these results, we have depicted all these objects in the HR diagram (Fig. 9) together with the evolutionary tracks for post-AGB stars and PNe central stars from Miller Bertolami (2016). To derive the masses we used the tracks with metallicity of , for stars with kpc. expected to be in the disc (upper panel), while for those with kpc we used the tracks. Using the luminosity threshold for post-AGB stars discussed before, the diagram allowed us to disclose bona fide post-AGB candidates from those that are not.
We would like to stress the fact that we are applying a quite restrictive selection threshold to sort out bona fide post-AGB candidates. It could certainly be the case that some other objects, located close to our luminosity threshold, could also be post-AGB but, given the uncertainties, do not fulfil our selection criteria, that we named as possible post-AGB stars.
It is worth noticing here that 12 of the objects that we classify as post-AGBs have been identified as possible, or confirmed, binary stars by Kluska et al. (2019). We checked the bibliography for other binary references, and found 5 additional binary stars. These stars together with two possible YSO binary stars, also according to Kluska et al. (2019), one supergiant star and one post-AGB unconfirmed candidate are identified with a flag in Table LABEL:table:candidates. It is worth noticing as well, that our SED classification is disc-type for all but one of those binary objects.
Figure 9 also shows two types of YSOs. All those objects with are suspected to be YSOs, but they could also be other types of evolved stars such as HB stars. There are also examples of well-known YSOs among objects with higher luminosities. The physical nature as YSOs of those luminous objects is more difficult to discern because their luminosities and temperatures overlap with those of post-AGB stars, and what we can do is resort to studying their locations in the Galaxy. If their positions and distances match those of star-forming regions, they can quite safely be classified as young objects.
We used the molecular clouds catalogue by Zucker et al. (2020), which gives coordinates and distances to a large number of those regions, and found that 5 of our objects are located within these clouds. Such objects are labeled as YSO in MoC (Molecular Clouds) in Figure 9 and Table A.4.
From those, 3 objects have luminosities above and 2 below that limit. The classification as YSOs of the remaining 8 objects with luminosities below is tentative, and consequently, we label them as possible YSOs in Fig. 9.
Figure 9 also shows 3 objects that we found to be Horizontal Branch stars: SDS2012 Ter8 38 is a blue Horizontal Branch star in the globular cluster Terzan 8, [SDS2012] NGC 6402 160 in NGC 6402 and BPS BS 16479-0009 is a field Horizontal Branch star candidate according to Beers et al. (1996). Finally, we found three objects above the luminosity upper limit expected for post-AGB stars that we classify as supergiant stars. We shall comment further on them in section 5.
3.6 Comparison with other classifications in the literature
We can compare our classification with that recently obtained by Aoki et al. (2022) for the 7 objects in common in both samples. 4 post-AGBs are identically classified by both of us, while 2 of our unconfirmed post-AGB stars were catalogued as post-AGB or as cool post-AGB by them and one of our possible YSOs was catalogued as a hot subdwarf by those authors.
When comparing our results with those in Kamath et al. (2022) we found 11 objects in common (those with good astrometric quality in that work, RUWE¡1.4). We can confirm their nature as bona fide post-AGB candidates for 10 of them while HD 107360 remains slightly underluminous for the post-AGB threshold. In general, the temperatures and luminosity values given by Kamath et al. (2022) are in agreement with our derived values, as it is illustrated in Table 1.
Object | ||||||
(mag) | (mag) | (K) | (K) | |||
HD 56126 | 2.0 | 1.33(*) | 7250 | 7485 | 3.94 | 3.74 |
HD 107369 | 0.3 | 0.22 | 7500 | 7533 | 2.99 | 2.96 |
HD 133656 | 1.0 | 0.90 | 8250 | 8238 | 3.74 | 3.72 |
HD 148743 | 0.6 | 0.34 | 6750 | 6728 | 4.51 | 4.41 |
HD 161796 | 0.75 | 0.40 | 6000 | 6139 | 3.88 | 3.76 |
HD 187885 | 1.8 | 1.74 | 8000 | 8239 | 3.89 | 3.85 |
HD 235858 | 2.8 | 2.73 | 5250 | 5325 | 3.94 | 3.75 |
IRAS 01259+6823 | 3.0 | 3.20 | 5500 | 5510 | 3.47 | 2.53 |
IRAS 12360-5740 | 2.7 | 3.10 | 7500 | 7273 | 3.88 | 3.80 |
V* LN Hya | 1.0 | 0.93 | 6250 | 6393 | 4.02 | 4.03 |
V* V1401 Aql | 1.0 | 1.24 | 6750 | 6985 | 3.47 | 3.55 |
(1): calculated by VOSA, (2): Kamath et al. (2022).
In Section 3.3 we compared our results with those obtained by Oudmaijer et al. (2022), as shown in Figures 5 and 7. For the Luminosity range expected for post-AGB stars, temperatures and luminosities agree with a dispersion that can be explained by methodological differences, as already discussed in that section. From 59 objects in common with Oudmaijer et al. (2022) sample, we have classified as post-AGB 44 of them, while we have catalogued 8 of them as unconfirmed candidates, 3 as possible YSOs, 2 as YSOs in molecular clouds and another 2 as supergiants.
Finally, we can summarise our results. Starting from the lists of post-AGB objects known or proposed as such in the literature, we have made a careful selection based on the quality of the astrometry in Gaia DR3 of these sources. We used updated dust extinction maps, 3D when available, for a more accurate derivation of luminosities. As a consequence, we were able to classify some of them as bona-fide post-AGB stars (69), supergiants (3), HB stars (3), YSOs in molecular clouds (5) and possible YSOs (9), while 29 objects remain as unconfirmed post-AGB candidates. In the following section, we describe the evolutionary properties of our sample of 69 post-AGB stars.
4 Sample of post-AGB stars
We focussed now on the sample of 69 objects whose luminosities allow us to confirm their evolutionary state as bona fide post-AGB candidates. By interpolating between the novel evolutionary models by Miller Bertolami (2016) for post-AGB stars, we estimate their progenitor mass (in the MS) and their evolutionary age in the post-AGB phase. Miller Bertolami (2016) provides tracks for metallicity values 0.01, 0.02, 0.001, and 0.0001. As we already mentioned, we used the tracks for as representative of the disc population, and for those objects belonging to the halo we used tracks.
Note that the objects classified as unconfirmed post-AGB candidates are located below the 1 track (0.9 for halo stars). It is assumed that these objects, if they are single-evolved stars, can only have initial masses slightly below 1 . Conversely, they could have their origin in binary evolution.
The progenitor mass distribution we obtained for the post-AGB stars in our sample is displayed in Fig. 10. About half of our stars (35) have masses below 1.5 in the MS, while only 5 of them have a mass higher than 3.5 . Although our study is limited in the number of objects and possibly comes from a biased selection, the resulting masses match the expected distribution. Parent stars with masses above 3.5 have too short lifetimes in the post-AGB phase (the crossing times for post-AGB and PNe phases in Miller Bertolami tracks are less than 400 yr) to populate this region. The mean value of the progenitor masses for the post-AGB sample is:
This mean mass value agrees with the value of obtained in González-Santamaría et al. (2021) for stars in the next evolutionary phase, as central stars of planetary nebulae (CSPNe).
The mean mass value for the post-AGB stars final sample resulted:
Regarding evolutionary ages, we obtained the distribution shown in Fig. 11. Around 36% of the stars have ages smaller than 2000 yrs. Moreover, almost all the objects in the sample show evolutionary ages below 10,000 yrs, with only 6 stars over that age. We have obtained the following evolutionary age mean value of the sample:
To be reminded is that in Miller Bertolami models, the beginning of the post-AGB phase is taken when, due to stellar winds, the mass of the external layer of the star drops below 1% of the star mass.
Individual mass and evolutionary age values can be found in columns 11 and 12 of Table 3, respectively.
5 Special objects
It deserves special mention that our study allows confirming the nature as post-AGB objects of three objects, lacking a previous firm classification in the catalogues, neither in the Torun catalogue as Likely, nor in Simbad or in the Suárez et al. catalogues as confirmed post-AGB.
CD-30 15464 – This object is catalogued as Possible post-AGB in the Torun catalogue, but as simple star in Simbad and it has no references in Suárez et al. (2006) catalogue. It is reported as a B1 spectral type star, while its effective temperature is quite high (22,000 K), still within the limits of a post-AGB star. It is located very far from the Sun, at a distance of 9.96 kpc, at the galactic centre direction () and slightly below the galactic disk (). Despite this location, the interstellar extinction is not very high in the direction of this star (), and its image777http://cdsportal.u-strasbg.fr/?target=CD-30%2015464 in PanSTARRS colours shows evidence of a circumstellar envelope.
HD 53300 – It is catalogued as A2 type star and it is located in the galactic disk (). Instead, Rao et al. (2012) derived an effective temperature of 7250 K for it, which corresponds to a F1 type star. This star is mentioned as a candidate post-AGB in the Simbad database and it has no references in the Torun or Suárez et al. (2006) catalogues, although Rao et al. (2012) classifies it as post-AGB star, while in Bhatt & Manoj (2000) it is identified as a Vega-like star.
HD 214539 – This star is catalogued as Possible post-AGB in Torun and as simple Star in Simbad, with a B8/9 spectral type. Kodaira & Philip (1984) obtained a temperature of 10,000 K and a gravity of log[g]=2 for this object. It is located at a distance of 1.39 kpc. By a visual inspection888http://cdsportal.u-strasbg.fr/?target=HD%20214539, also in this case, a circumstellar envelope typical of post-AGB stars can be observed.
It is also interesting to analyse the three highly luminous objects () that we have catalogued as supergiant stars.
BD-02 4931 – This object is catalogued as post-AGB candidate star in the Simbad database, while in Parthasarathy et al. (2000a) it is classified as a B1 type giant. Assuming a temperature of 16,000 K from its spectral type (the same value as the one obtained by fitting its SED by Gielen et al. 2011) we obtain a luminosity of , higher than the predictions for post-AGB stars.
HD 179821 – This star is catalogued as Likely post-AGB in the Torun catalogue and as a post-AGB star in the Simbad database and in Suárez et al. catalogue. However, we have obtained a high luminosity of (very similar to that obtained by Wood et al. (1983), ) and low surface gravity of for this object, that leads us to classify it as a supergiant star. Moreover, in Şahin et al. (2016), they catalogued it as a likely massive post-red supergiant star.
V* V1027 Cyg – This object is also classified as Likely post-AGB in the Torun catalogue and as post-AGB in Simbad database. But, in this case, we obtained stellar parameters typical of supergiant stars, such as a high luminosity of and a very low surface gravity of . Furthermore, in the Winfrey et al. (1994) catalogue, they classified it as a G7 supergiant star, while it is catalogued as a G8-K3 type supergiant in a more recent study by Arkhipova et al. (2016).
6 Conclusions
From a sample of 118 post-AGB stars candidates selected from the literature and having filtered out those with Gaia DR3 not accurate enough astrometry (and hence unreliable distances), we have estimated their luminosity values which allow us to classify them as bona fide post-AGB candidates (69) or as objects in a different evolutionary phase, such as YSOs (5), possible YSOs (9), supergiant stars (3) and HB stars (3), while 29 stars remain as unconfirmed post-AGB candidates.
Using the spanish VOSA service, we have fitted the SEDs of each star and simultaneously obtained its effective temperature and luminosity. This allows us to plot the post-AGB candidates in an HR diagram, and by using Miller Bertolami (2016) evolutionary tracks in the post-AGB phase, to derive their masses and ages. We found that our 69-object sample includes mainly stars with progenitor masses between 1 and 2.5 values that agree with the type of post-AGB stars that statistically could be found in a small sample like ours. The mass mean value of the sample is also in agreement with that expected for stars in the planetary nebula phase (the next evolutionary phase), according to González-Santamaría et al. (2021).
Our study allows us to confirm the nature as post-AGB stars of several objects previously unconfirmed as such. It is important to note that although the Gaia DR3 catalogue contains statistically valuable information for an enormous number of stars, many of the available parallaxes contain errors too large to allow a correct estimation of the absolute magnitude or luminosity of the stars. We chose to perform rather restrictive filtering of the astrometric quality of Gaia measurements (parallax and distance errors, RUWE value) as well as of the distance inferred from the parallax measurement by a Bayesian model of galactic stellar distributions.
Behind this filtering is the idea of working with a small set of objects, candidates to be in the post-AGB phase, all of them with precise distances in Gaia, which therefore will mostly be individual objects for which a luminosity can be calculated with great confidence. For this reason, our methodology was based on selecting only a subset of stars that are firm candidates to be in the post-AGB phase and exhibit Gaia DR3 parallaxes (and inferred distances) with errors below 30%.
Also to be noticed is that we worked only with those candidates with interstellar extinction values from Stassun 3D dust maps, which allowed us to initially constrain interstellar extinction. Total extinction, including both interstellar and circumstellar extinction, was derived simultaneously with the luminosity value from the SED fitting procedure. Effective temperatures were taken from the literature, using spectroscopically determined temperatures when they were available.
We discuss in detail our results in comparison with other similar studies about post-AGBs, carried out recently. About 25% of the stars in our sample resulted underluminous to be confirmed as post-AGB stars under our working hypotheses, while 12% are found to be possible YSOs, either located in Molecular Clouds (5) or candidates (9). Other objects such as HB stars (3) or supergiant stars (3) were also included in the original compilation.
Regarding binarity, although our initial filtering would rule out most of the binary objects, we found that 18 objects out of our sample of 69 post-AGB have SEDs that can be classified as disk-type and, therefore, might correspond to binary objects. Searching the literature for binarity, we found that 17 objects among them have been identified as such either by Kluska et al. 2022 (12 objects) or by other authors (see table A.1 for references). This brings us to some 18 possible or confirmed binaries within our sample, about 26%.
We think our results provide an interesting framework to pursue further insight into the study of the post-AGB phase. In particular, given our well characterised sample of 69 objects, a way is open to complement the study on the unconfirmed 29) candidates proposed. A follow-up analysis of their properties, including spectroscopy when possible, would be desirable.
Acknowledgements.
This work has made use of data from the European Space Agency (ESA) Gaia mission, processed by the Gaia Data Processing and Analysis Consortium (DPAC). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement. This research has made use of the Simbad database and the Aladin sky atlas, operated at CDS, Strasbourg, France. The authors have also made use of the VOSA software, developed under the Spanish Virtual Observatory project supported by the Spanish MINECO through grant PID2020-112949GB-I00, and partially funded by the European Union’s Seventh Framework Programme (FP7-SPACE-2013-1) for research, technological development and demonstration under grant agreement no. 60674. Funding from Spanish Ministry project PID2021-122842OB-C22, Xunta de Galicia ED431B 2021/36 and PDC2021-121059-C22 is acknowledged by the authors. We also acknowledge support from CIGUS-CITIC, funded by Xunta de Galicia and the European Union (FEDER Galicia 2014-2020 Program) through grant ED431G 2019/01. E.V. acknowledges support from the ’DISCOBOLO’ funded by the Spanish Ministerio de Ciencia, Innovación y Universidades under grant PID2021-127289NB-I00. M.M. and E.V. acknowledge support from the cooperation agreement between the IAC and the Fundación Jesús Serra for visiting grants. AM acknowledges support from the ACIISI, Gobierno de Canarias and the European Regional Development Fund (ERDF) under grant with reference PROID2020010051 as well as from the State Research Agency (AEI) of the Spanish Ministry of Science and Innovation (MICINN) under grant PID2020-115758GB-I00.References
- Aoki et al. (2022) Aoki, W., Matsuno, T., & Parthasarathy, M. 2022, PASJ, 74, 1368
- Arentsen et al. (2019) Arentsen, A., Prugniel, P., Gonneau, A., et al. 2019, A&A, 627, A138
- Arkhipova et al. (2016) Arkhipova, V. P., Taranova, O. G., Ikonnikova, N. P., et al. 2016, Astronomy Letters, 42, 756
- Bailer-Jones et al. (2021) Bailer-Jones, C. A. L., Rybizki, J., Fouesneau, M., Demleitner, M., & Andrae, R. 2021, AJ, 161, 147
- Bayo et al. (2008) Bayo, A., Rodrigo, C., Barrado Y Navascués, D., et al. 2008, A&A, 492, 277
- Beers et al. (1996) Beers, T. C., Wilhelm, R., Doinidis, S. P., & Mattson, C. J. 1996, ApJS, 103, 433
- Bhatt & Manoj (2000) Bhatt, H. C. & Manoj, P. 2000, A&A, 362, 978
- Bloecker (1995) Bloecker, T. 1995, A&A, 299, 755
- Bonnarel et al. (2000) Bonnarel, F., Fernique, P., Bienaymé, O., et al. 2000, A&AS, 143, 33
- Buder et al. (2021) Buder, S., Sharma, S., Kos, J., et al. 2021, MNRAS, 506, 150
- Castelli & Kurucz (2003) Castelli, F. & Kurucz, R. L. 2003, in Modelling of Stellar Atmospheres, ed. N. Piskunov, W. W. Weiss, & D. F. Gray, Vol. 210, A20
- Corporaal et al. (2023) Corporaal, A., Kluska, J., Van Winckel, H., et al. 2023, A&A, 674, A151
- Şahin (2018) Şahin, T. 2018, Astrophysical Bulletin, 73, 211
- Şahin et al. (2016) Şahin, T., Lambert, D. L., Klochkova, V. G., & Panchuk, V. E. 2016, MNRAS, 461, 4071
- de Ruyter et al. (2005) de Ruyter, S., van Winckel, H., Dominik, C., Waters, L. B. F. M., & Dejonghe, H. 2005, A&A, 435, 161
- Doroshenko et al. (2016) Doroshenko, V., Pühlhofer, G., Kavanagh, P., et al. 2016, MNRAS, 458, 2565
- Drilling et al. (2013) Drilling, J. S., Jeffery, C. S., Heber, U., Moehler, S., & Napiwotzki, R. 2013, A&A, 551, A31
- Dubus et al. (2018) Dubus, G., Otulakowska-Hypka, M., & Lasota, J.-P. 2018, A&A, 617, A26
- Fabricius et al. (2021) Fabricius, C., Luri, X., Arenou, F., et al. 2021, A&A, 649, A5
- Firnstein & Przybilla (2012) Firnstein, M. & Przybilla, N. 2012, A&A, 543, A80
- Fitzpatrick (1999) Fitzpatrick, E. L. 1999, PASP, 111, 63
- Gallardo Cava et al. (2023) Gallardo Cava, I., Alcolea, J., Bujarrabal, V., Gómez-Garrido, M., & Castro-Carrizo, A. 2023, A&A, 671, A80
- Garcia-Lario et al. (1997) Garcia-Lario, P., Manchado, A., Pych, W., & Pottasch, S. R. 1997, A&AS, 126, 479
- Gielen et al. (2011) Gielen, C., Bouwman, J., van Winckel, H., et al. 2011, A&A, 533, A99
- Gonzalez & Wallerstein (1992) Gonzalez, G. & Wallerstein, G. 1992, MNRAS, 254, 343
- González-Santamaría et al. (2021) González-Santamaría, I., Manteiga, M., Manchado, A., et al. 2021, A&A, 656, A51
- Hardorp et al. (1965) Hardorp, J., Theile, I., & Voigt, H. H. 1965, Hamburger Sternw. Warner & Swasey Obs., C05, 0
- Henize (1976) Henize, K. G. 1976, ApJS, 30, 491
- Herrero et al. (2020) Herrero, A., Parthasarathy, M., Simón-Díaz, S., et al. 2020, MNRAS, 494, 2117
- Hunger & Kaufmann (1973) Hunger, K. & Kaufmann, J. P. 1973, A&A, 25, 261
- Ikonnikova et al. (2020) Ikonnikova, N. P., Parthasarathy, M., Dodin, A. V., Hubrig, S., & Sarkar, G. 2020, MNRAS, 491, 4829
- Indebetouw et al. (2005) Indebetouw, R., Mathis, J. S., Babler, B. L., et al. 2005, ApJ, 619, 931
- Jeffery (1993) Jeffery, C. S. 1993, A&A, 279, 188
- Jeffery et al. (1998) Jeffery, C. S., Hamill, P. J., Harrison, P. M., & Jeffers, S. V. 1998, A&A, 340, 476
- Kamath et al. (2022) Kamath, D., Van Winckel, H., Ventura, P., et al. 2022, ApJ, 927, L13
- Kamath et al. (2015) Kamath, D., Wood, P. R., & Van Winckel, H. 2015, MNRAS, 454, 1468
- Klochkova (2014) Klochkova, V. G. 2014, Astrophysical Bulletin, 69, 279
- Klochkova & Chentsov (2018) Klochkova, V. G. & Chentsov, E. L. 2018, Astronomy Reports, 62, 19
- Klochkova et al. (2018) Klochkova, V. G., Sendzikas, E. G., & Chentsov, E. L. 2018, Astrophysical Bulletin, 73, 52
- Kluska et al. (2022) Kluska, J., Van Winckel, H., Coppée, Q., et al. 2022, A&A, 658, A36
- Kluska et al. (2019) Kluska, J., Van Winckel, H., Hillen, M., et al. 2019, A&A, 631, A108
- Kodaira & Philip (1984) Kodaira, K. & Philip, A. G. D. 1984, ApJ, 278, 208
- Lada et al. (2006) Lada, C. J., Muench, A. A., Luhman, K. L., et al. 2006, AJ, 131, 1574
- Lindegren et al. (2021) Lindegren, L., Bastian, U., Biermann, M., et al. 2021, A&A, 649, A4
- Lindegren et al. (2018) Lindegren, L., Hernández, J., Bombrun, A., et al. 2018, A&A, 616, A2
- Luck (2014) Luck, R. E. 2014, AJ, 147, 137
- Maas et al. (2005) Maas, T., Van Winckel, H., & Lloyd Evans, T. 2005, A&A, 429, 297
- Mello et al. (2012) Mello, D. R. C., Daflon, S., Pereira, C. B., & Hubeny, I. 2012, A&A, 543, A11
- Miller Bertolami (2016) Miller Bertolami, M. M. 2016, A&A, 588, A25
- Mooney et al. (2004) Mooney, C. J., Rolleston, W. R. J., Keenan, F. P., et al. 2004, A&A, 419, 1123
- Napiwotzki et al. (1994) Napiwotzki, R., Heber, U., & Koeppen, J. 1994, A&A, 292, 239
- Oomen et al. (2018) Oomen, G.-M., Van Winckel, H., Pols, O., et al. 2018, A&A, 620, A85
- Oudmaijer et al. (2022) Oudmaijer, R. D., Jones, E. R. M., & Vioque, M. 2022, MNRAS, 516, L61
- Parthasarathy et al. (2022) Parthasarathy, M., Kounkel, M., & Stassun, K. G. 2022, Research Notes of the American Astronomical Society, 6, 171
- Parthasarathy et al. (2020) Parthasarathy, M., Matsuno, T., & Aoki, W. 2020, PASJ, 72, 99
- Parthasarathy et al. (2000a) Parthasarathy, M., Sivarani, T., Garcia-Lario, P., & Manchado, A. 2000a, in American Astronomical Society Meeting Abstracts, Vol. 197, American Astronomical Society Meeting Abstracts, 60.05
- Parthasarathy et al. (2000b) Parthasarathy, M., Vijapurkar, J., & Drilling, J. S. 2000b, A&AS, 145, 269
- Quin & Lamers (1992) Quin, D. A. & Lamers, H. J. G. L. M. 1992, A&A, 260, 261
- Raman et al. (2017) Raman, V. V., Anandarao, B. G., Janardhan, P., & Pandey, R. 2017, MNRAS, 470, 1593
- Rao et al. (2012) Rao, S. S., Giridhar, S., & Lambert, D. L. 2012, MNRAS, 419, 1254
- Reyniers & Van Winckel (2001) Reyniers, M. & Van Winckel, H. 2001, A&A, 365, 465
- Stassun (2019) Stassun, K. G. 2019, VizieR Online Data Catalog, IV/38
- Steinmetz et al. (2020) Steinmetz, M., Guiglion, G., McMillan, P. J., et al. 2020, AJ, 160, 83
- Suárez et al. (2006) Suárez, O., García-Lario, P., Manchado, A., et al. 2006, A&A, 458, 173
- Szczerba et al. (2007) Szczerba, R., Siódmiak, N., Stasińska, G., & Borkowski, J. 2007, A&A, 469, 799
- Szczerba et al. (2012) Szczerba, R., Siódmiak, N., Stasińska, G., et al. 2012, IAU Symposium, 283, 506
- Taylor (2005) Taylor, M. B. 2005, in Astronomical Society of the Pacific Conference Series, Vol. 347, Astronomical Data Analysis Software and Systems XIV, ed. P. Shopbell, M. Britton, & R. Ebert, 29
- Ting et al. (2019) Ting, Y.-S., Conroy, C., Rix, H.-W., & Cargile, P. 2019, ApJ, 879, 69
- van Aarle et al. (2011) van Aarle, E., van Winckel, H., Lloyd Evans, T., et al. 2011, VizieR Online Data Catalog, J/A+A/530/A90
- Van Winckel (1997) Van Winckel, H. 1997, A&A, 319, 561
- Van Winckel et al. (2014) Van Winckel, H., Jorissen, A., Exter, K., et al. 2014, A&A, 563, L10
- Vassiliadis & Wood (1993) Vassiliadis, E. & Wood, P. R. 1993, ApJ, 413, 641
- Venn et al. (1998) Venn, K. A., Smartt, S. J., Lennon, D. J., & Dufton, P. L. 1998, A&A, 334, 987
- Vickers et al. (2015) Vickers, S. B., Frew, D. J., Parker, Q. A., & Bojičić, I. S. 2015, MNRAS, 447, 1673
- Villaver et al. (2002) Villaver, E., Manchado, A., & García-Segura, G. 2002, ApJ, 581, 1204
- Waelkens et al. (1991) Waelkens, C., Van Winckel, H., Bogaert, E., & Trams, N. R. 1991, A&A, 251, 495
- Wallerstein (1958) Wallerstein, G. 1958, ApJ, 127, 583
- Winfrey et al. (1994) Winfrey, S., Barnbaum, C., Morris, M., & Omont, A. 1994, in American Astronomical Society Meeting Abstracts, Vol. 185, American Astronomical Society Meeting Abstracts, 45.15
- Wood et al. (1983) Wood, P. R., Bessell, M. S., & Fox, M. W. 1983, ApJ, 272, 99
- Zucker et al. (2020) Zucker, C., Speagle, J. S., Schlafly, E. F., et al. 2020, A&A, 633, A51
Appendix A Data Tables
Num | Simbad Name | Gaia DR3 ID | RA | Dec | G mag | Spectral Type | Reference | Flag | |
(º) | (º) | ||||||||
1 | * 42 Cyg | 2056972679739120000 | 307.335 | 36.4547 | 5.74 | 1.52 | A2Iab-Ib | 1 | - |
2 | * 89 Her | 4582795323914832000 | 268.8549 | 26.05 | 5.35 | 0.17 | F2Ibp | 1/2 | B(1,2,3) |
3 | [SDS2012] NGC 6284 116 | 4112726164275562496 | 256.1875 | -24.55 | 16.93 | - | - | 1 | - |
4 | [SDS2012] NGC 6402 160 | 4368932547017571584 | 264.1583 | -3.3867 | 16.85 | 1.64 | - | 1 | - |
5 | [SDS2012] Ter 8 38 | 6741750932641613952 | 295.42 | -34.0664 | 15.09 | 0.41 | - | 1 | - |
6 | 2MASS J00235767-7205296 | 4689637789379454848 | 5.9903 | -72.0916 | 11.09 | 0.09 | - | 1 | - |
7 | 2MASS J01302276-7303339 | 4686479648370453760 | 22.5949 | -73.0594 | 10.63 | 0.12 | B | 1/2 | - |
8 | 2MASS J05241036-2429206 | 2957941232276476800 | 81.0432 | -24.4891 | 12.09 | 0.09 | - | 1 | - |
9 | 2MASS J06544616-1048325 | 3049274119844532224 | 103.6924 | -10.8091 | 12.05 | 2.03 | OB+ | 1 | R(1) |
10 | 2MASS J13272898-4722472 | 6083708479176016128 | 201.8708 | -47.3798 | 13.09 | 0.37 | - | 1/2 | R(1) |
11 | 2MASS J14034398-6937097 | 5846979777414443264 | 210.9333 | -69.6194 | 16.39 | 0.79 | - | 1 | - |
12 | 2MASS J16570924-0404243 | 4365635249084745600 | 254.2886 | -4.0734 | 13.49 | 0.79 | B | 1 | R(1) |
13 | 2MASS J17390218-4500388 | 5955201232284272384 | 264.7591 | -45.0108 | 13.01 | 1.15 | - | 1/2 | R(1) |
14 | 2MASS J17442550-1937537 | 4119884023727270272 | 266.1063 | -19.6316 | 12.43 | 1.67 | OB+ | 1 | - |
15 | 2MASS J18224265-3014383 | 4046476465531783424 | 275.6777 | -30.244 | 11.8 | 0.39 | B7Ib | 1/2 | - |
16 | 2MASS J18530579-0842378 | 4203848980711226112 | 283.2741 | -8.7105 | 13.14 | 1.14 | - | 1/2 | - |
17 | BD+32 2754 | 1324742534573959424 | 249.0487 | 32.4893 | 9.47 | 0.07 | F8 | 2 | - |
18 | BD+33 2642 | 1369896865785991424 | 237.9995 | 32.9484 | 10.79 | 0.08 | B2 IVp | 1/2 | B(4) |
19 | BD+48 1220 | 255225480926107392 | 76.9595 | 48.4026 | 9.53 | 0.89 | A4Ia | 1/2 | - |
20 | BD-02 4931 | 4213102543594938880 | 289.5947 | -2.703 | 10.42 | 1.68 | B1III | 1 | B(5) |
21 | BD-13 5550 | 6879196723703009920 | 300.4576 | -12.6883 | 11.34 | 0.44 | B1Iae | 1/2 | - |
22 | BPS BS 16479-0009 | 3939536182204010880 | 198.4999 | 18.5253 | 13.61 | 0.06 | - | 1 | - |
23 | CD-24 13065 | 4113478337639987200 | 256.0433 | -24.4661 | 11.08 | 0.47 | B8 | 1 | - |
24 | CD-30 15464 | 4049379725984965120 | 274.002 | -30.7565 | 11.9 | 0.75 | B1Ib | 2 | - |
25 | CD-42 8141 | 6135778223095320960 | 197.1931 | -43.4642 | 10.45 | 0.34 | B2I | 1 | - |
26 | CD-46 11775 | 5948818331093816448 | 265.6416 | -46.9802 | 11.17 | 0.61 | OB+ | 1 | - |
27 | CD-48 11445 | 5938738764416910848 | 256.9027 | -48.319 | 10.5 | 2.24 | G2p(R) | 1 | B(1,5) |
28 | CD-49 8217 | 6093466301247487488 | 207.3233 | -50.3793 | 10.82 | 0.61 | B2I | 1 | R(1) |
29 | CD-49 11554 | 5946845601071213696 | 263.7604 | -49.4407 | 10.93 | 0.61 | B3Ie | 1/2 | - |
30 | CD-53 5736 | 5893945588395282304 | 223.1197 | -54.2952 | 10.87 | 2.53 | A0Ie | 1/2/3 | - |
31 | CD-54 5573 | 5896479309853592448 | 212.6621 | -55.0075 | 10.29 | 1.47 | A3I | 1/2 | - |
32 | CD-54 6746 | 5932016212933920384 | 246.1642 | -54.6357 | 9.45 | 1.11 | B8Iab | 1/2 | - |
33 | CD-55 5174 | 6063703586653222144 | 202.4626 | -56.1149 | 10.71 | 1.26 | B1Iae | 1/2/3 | - |
34 | CD-59 6142 | 5831295999979910656 | 246.2609 | -60.059 | 9.98 | 0.59 | A3Ie | 1/2/3 | - |
35 | Cl* NGC 6779 SAW V6 | 2039259886717168896 | 289.1491 | 30.1941 | 12.53 | 0.56 | kF5hF8 | 1 | - |
36 | EM* GGR 44 | 2005246464463628800 | 331.0513 | 53.0671 | 12.43 | 1.11 | B1I | 1/2 | R(1) |
37 | EM* StHA 161 | 2049984454412871296 | 290.4804 | 35.0486 | 11.31 | 0.27 | Fe | 1/2/3 | - |
38 | EM* VES 351 | 2168803045330976768 | 314.7316 | 49.5203 | 11.18 | 2.06 | F3Ie | 1/2/3 | - |
39 | HD 53300 | 3101342596792542464 | 106.0812 | -5.3054 | 7.94 | 0.98 | A2II | 1 | - |
40 | HD 56126 | 3156171118495247360 | 109.0427 | 9.9967 | 8.1 | 0.08 | F0/5Ia | 1/2/3 | - |
41 | HD 93662 | 5351069693654349952 | 161.9101 | -57.4674 | 5.66 | 0.37 | K5 | 1/2 | B(1,2) |
42 | HD 101584 | 5343168568718268800 | 175.245 | -55.5738 | 6.92 | 0.82 | F0Iape | 1/2 | B(1,2,8) |
43 | HD 105262 | 3920735495441657728 | 181.7951 | 12.9855 | 7.07 | 0.05 | B9 Ib | 1/2 | B(8) |
44 | HD 107369 | 3469106382752903168 | 185.1873 | -32.5573 | 9.54 | 0.21 | A2II/III | 1/2 | - |
45 | HD 108015 | 6130448958959242240 | 186.2229 | -47.1521 | 7.86 | 0.31 | F3/5Ib/II | 1/2 | B(1,2) |
46 | HD 116745 | 6083719439934104832 | 201.6097 | -47.2743 | 10.68 | 0.36 | F0Ibp | 1/2 | - |
47 | HD 133656 | 5903310335089068416 | 226.8643 | -48.2983 | 7.49 | 0.74 | A1/A2Ib/II | 1/2/3 | - |
48 | HD 144941 | 6042510190769087744 | 242.3523 | -27.2273 | 10.09 | 0.72 | B8 | 1 | - |
49 | HD 148743 | 4351018375858237952 | 247.6251 | -7.5145 | 6.37 | 0.67 | A7Ib | 1/2 | - |
50 | HD 157350 | 4122877783340594176 | 260.8558 | -17.971 | 8.56 | 0.1 | A2III/IV | 1 | - |
51 | HD 161796 | 1367102319545324288 | 266.2311 | 50.0443 | 7.2 | 0.11 | F3Ib | 1/2/3 | - |
52 | HD 167402 | 4049624646596488576 | 274.0779 | -30.1249 | 8.94 | 0.73 | O9.5/B0Ib/II | 1 | R(1) |
53 | HD 172324 | 2096072103492979584 | 279.4949 | 37.4349 | 8.16 | 0.11 | A0Iabe | 1/2 | - |
54 | HD 172481 | 4072427555640528000 | 280.404 | -27.9503 | 8.84 | 0.65 | F2/3Ia | 1/2 | B(1,2) |
55 | HD 177566 | 6715619076008049792 | 286.7829 | -41.7211 | 10.14 | 0.26 | B6Ib | 1/2 | R(1) |
56 | HD 179821 | 4264026012336768000 | 288.4942 | 0.1255 | 7.55 | 1.25 | 1/2/3 | - | |
57 | HD 186438 | 2049034819957965312 | 295.7205 | 37.6782 | 7.83 | 0.26 | F3Ib | 1/2 | - |
58 | HD 187885 | 6871175064823382912 | 298.2196 | -17.0307 | 8.48 | 0.62 | F0Ie | 1/2/3 | - |
59 | HD 214539 | 6385794694664872320 | 340.1999 | -67.6886 | 7.2 | 0.09 | B8/9I | 2 | - |
60 | HD 235858 | 2006425553228658816 | 337.2933 | 54.8517 | 8.2 | 0.48 | G5Ia | 1/2 | - |
61 | HD 246299 | 3336558507975208448 | 85.2377 | 10.2403 | 10.27 | 0.58 | G2I | 1/2/3 | - |
62 | HD 306753 | 5335709477519159936 | 174.4288 | -60.8976 | 12.36 | 2.01 | A0 | 1/2 | - |
63 | IRAS 01005+7910 | 565507868441719424 | 16.1896 | 79.4462 | 10.96 | 0.42 | B2Iab | 1/2/3 | - |
64 | IRAS 01259+6823 | 532078488712794624 | 22.3892 | 68.6547 | 11.81 | 2.74 | F5Ie | 2 | - |
65 | IRAS 02528+4350 | 433515788197481984 | 44.0473 | 44.0478 | 10.75 | 0.34 | A0e | 1/2 | - |
66 | IRAS 07227-1320 | 3032030620730261376 | 111.2628 | -13.4389 | 11.6 | 0.66 | M1I | 2 | - |
67 | IRAS 07582-4059 | 5534265613756612224 | 119.9905 | -41.1231 | 14.55 | 2.5 | - | 2/3 | R(2) |
68 | IRAS 08242-3828 | 5540178478053582592 | 126.5158 | -38.6465 | 12.03 | - | - | 1/2 | - |
69 | IRAS 08275-6206 | 5277809440015969792 | 127.1014 | -62.2724 | 10.8 | 0.69 | - | 1/2 | - |
70 | IRAS 08351-4634 | 5521628033275348480 | 129.19 | -46.7469 | 17.13 | 1.64 | - | 1/2 | R(3) |
71 | IRAS 09370-4826 | 5409357863031443840 | 144.728 | -48.6731 | 13.75 | 0.41 | - | 2 | R(3) |
72 | IRAS 11387-6113 | 5335675087769798272 | 175.2863 | -61.5048 | 11.4 | 1.41 | A3Ie | 1/2/3 | - |
73 | IRAS 11531-6111 | 5335102207846402176 | 178.9084 | -61.4713 | 14.49 | 2.65 | - | 2 | - |
74 | IRAS 12145-5834 | 6071416385848395008 | 184.3171 | -58.8582 | 15.09 | 2.55 | B8Ie | 2 | R(4) |
75 | IRAS 12360-5740 | 6060828565581083264 | 189.7213 | -57.9422 | 12.08 | 1.62 | - | 1/2 | - |
76 | IRAS 13110-6629 | 5857811238294426752 | 198.6128 | -66.7594 | 10.54 | - | - | 1/2 | - |
77 | IRAS 13356-6249 | 5865398796206273152 | 204.7763 | -63.079 | 15.8 | - | - | 1/2 | - |
78 | IRAS 13421-6125 | 5865808020691983104 | 206.392 | -61.6677 | 16.18 | 5.47 | - | 2 | R(3) |
79 | IRAS 14527-6204 | 5874676853324862720 | 224.1857 | -62.2815 | 11.08 | 1.74 | - | 2 | R(5) |
80 | IRAS 15066-5532 | 5886573569080505216 | 227.6111 | -55.7367 | 14.57 | 5.55 | - | 1/2 | - |
81 | IRAS 16086-5255 | 5933063252888145920 | 243.1269 | -53.0528 | 13.15 | 3.28 | - | 1/2 | - |
82 | IRAS 16115-5044 | 5935061172876722176 | 243.8248 | -50.8721 | 14.83 | - | - | 2 | - |
83 | IRAS 16476-1122 | 4334241408966611328 | 252.6012 | -11.466 | 11.09 | 1.7 | M1I | 2 | - |
84 | IRAS 16494-3930 | 5969973999973524224 | 253.233 | -39.5818 | 16.73 | 1.83 | G2I | 1 | - |
85 | IRAS 16594-4656 | 5963059480546004608 | 255.792 | -47.0077 | 14.6 | - | - | 1/2 | - |
86 | IRAS 17208-3859 | 5972489407685976320 | 261.0812 | -39.0294 | 15.46 | 2.45 | - | 2 | - |
87 | IRAS 17223-2659 | 4109553493474085504 | 261.361 | -27.0337 | 15.14 | 4.36 | - | 2 | - |
88 | IRAS 17287-3443 | 5975119332093959552 | 263.0201 | -34.7591 | 12.92 | 4.53 | - | 2 | B(6) |
89 | IRAS 17310-3432 | 4053542580087893376 | 263.585 | -34.5815 | 15.77 | 3.16 | - | 2 | R(1) |
90 | IRAS 17332-2215 | 4117592469707529856 | 264.0712 | -22.2889 | 15.22 | 2.27 | - | 2 | - |
91 | IRAS 17364-1238 | 4161796857143755264 | 264.8205 | -12.6749 | 12.8 | 1.76 | - | 2 | - |
92 | IRAS 17433-1750 | 4120632688077368192 | 266.5659 | -17.8628 | 13.33 | 1.64 | M2I | 1/2 | R(3) |
93 | IRAS 17543-3102 | 4044070253255414656 | 269.39 | -31.051 | 14.84 | - | - | 2 | - |
94 | IRAS 17579-3121 | 4043901443659288448 | 270.3057 | -31.3657 | 11.2 | - | - | 1/3 | - |
95 | IRAS 17581-2926 | 4062288993280721792 | 270.314 | -29.4441 | 11 | 1.59 | - | 2 | R(5) |
96 | IRAS 18084-1737 | 4095941436385621888 | 272.8678 | -17.611 | 15.71 | 1.07 | G3I | 1 | R(4) |
97 | IRAS 18113-2503 | 4065347387968755328 | 273.6136 | -25.0501 | 15.07 | 3.01 | - | 1/2 | R(3) |
98 | IRAS 18158-3445 | 4044520262548753152 | 274.8057 | -34.7417 | 12.45 | 0.42 | F6 | 1 | B(7) |
99 | IRAS 18435-0052 | 4260301176152524032 | 281.5326 | -0.8114 | 10.93 | 3.28 | B2II | 2 | - |
100 | IRAS 19075+0432 | 4293369057089082112 | 287.4997 | 4.619 | 14.44 | 4.79 | - | 1/2 | - |
101 | IRAS 19225+3013 | 2038872686817523072 | 291.1122 | 30.3241 | 12.24 | 0.67 | M2II | 2 | - |
102 | IRAS 19454+2920 | 2031794791233840128 | 296.8534 | 29.4697 | 15.29 | 4.43 | C-rich | 1/2 | - |
103 | IRAS 20094+3721 | 2060806470651334912 | 302.82 | 37.5145 | 10.63 | 1.24 | - | 1/2 | B(7) |
104 | IRAS 20174+3222 | 2054521833963867008 | 304.8659 | 32.5376 | 15.26 | - | - | 1/2 | - |
105 | IRAS 20244+3509 | 2056435602670418688 | 306.6049 | 35.3204 | 13.97 | 5.34 | - | 1/2 | - |
106 | IRAS 20259+4206 | 2068126263125039488 | 306.9261 | 42.2789 | 13.54 | 1.32 | - | 2 | - |
107 | IRAS 20490+5934 | 2193902559325301760 | 312.5566 | 59.7642 | 10.37 | 0.63 | A3e | 1/2 | - |
108 | IRAS 21289+5815 | 2179471159976791168 | 322.5951 | 58.4811 | 14.46 | 1.35 | A2Ie | 1/2 | - |
109 | IRAS 21525+5643 | 2198987491374918528 | 328.5631 | 56.957 | 16.69 | 2.53 | - | 2 | - |
110 | LB 3193 | 4715635535640762240 | 19.7214 | -61.9281 | 12.66 | 0.05 | - | 1/2 | - |
111 | LS IV -04 1 | 4365451214021224320 | 254.1155 | -4.7899 | 12 | 0.86 | B | 1/2 | - |
112 | LS IV -15 3 | 4136944866387751552 | 260.7996 | -15.6209 | 11.76 | 1.21 | A0Ie | 1/2/3 | - |
113 | LS 4331 | 4124125282361429504 | 265.2502 | -16.3035 | 13.08 | 1.39 | B1Ibe | 1/2 | - |
114 | LS 5112 | 4099619470274753408 | 280.2026 | -17.0773 | 11.88 | 1.32 | B1IIIep | 1/2 | - |
115 | LS II +34 26 | 1869422453048750336 | 312.0693 | 34.4567 | 11.05 | 0.62 | B1.5Ia | 2 | R(4) |
116 | LSE 63 | 6736747708089687936 | 280.0917 | -31.9469 | 12.04 | 0.42 | B1Iabe | 1/2 | - |
117 | NGC 6254 1035 | 4365635279142583424 | 254.299 | -4.0666 | 11.43 | 0.84 | G0e | 1 | - |
118 | OH 15.7 +0.8 | 4146237904302941440 | 274.1057 | -14.921 | 16.29 | 1.06 | - | 2 | R(3) |
119 | OH 17.7 -2.0 | 4104128533107792512 | 277.6288 | -14.4793 | 13.65 | 1.55 | - | 2 | R(3) |
120 | OH 345.05 -1.86 | 5965644393737729664 | 258.0907 | -42.4193 | 15.61 | 2.79 | - | 2 | R(3) |
121 | PG 1704+222 | 4568163710366782848 | 256.6924 | 22.0978 | 12.69 | 0.23 | sdB3IHe8 | 2 | - |
122 | PHL 1580 | 6831062200578042624 | 322.6052 | -19.3762 | 12.18 | 0.1 | B2 | 1/2 | R(1) |
123 | PN G038.7+01.5 | 4282499452616310912 | 284.0942 | 5.8833 | 12.5 | 2.91 | - | 2 | R(4) |
124 | PN PM 1-243 | 4104509518206051968 | 278.7396 | -13.9802 | 14.43 | 1.44 | - | 2 | R(4) |
125 | RAFGL 6945S | 4093773852321518976 | 272.8321 | -21.9166 | 13.99 | 1.7 | - | 2 | R(3) |
126 | SS 441 | 4210278482327706496 | 294.073 | -3.8903 | 13.06 | - | - | 1 | - |
127 | V* AD Aql | 4203801018864386944 | 284.7862 | -8.1706 | 11.28 | 0.78 | kF1hF5cnG5Ib | 1 | B(7)/V(1) |
128 | V* AU Vul | 1836195688380634368 | 304.5245 | 27.7343 | 9.98 | 1.57 | F3Ie | 1/2/3 | B(7)/V(1) |
129 | V* BZ Pyx | 5636099047820103424 | 137.0422 | -28.3196 | 10.91 | 0.35 | F6Ia | 1 | B(7)/V(1) |
130 | V* CE Vir | 3658327596544582400 | 207.3213 | -1.9291 | 8.31 | 0.15 | G8III | 1 | B(1) |
131 | V* EQ Cas | 1993916856117284864 | 358.222 | 55.0136 | 11.3 | - | - | 1 | - |
132 | V* HP Lyr | 2101097215232231808 | 290.4128 | 39.9356 | 10.41 | 0.33 | A3Ia/Iab | 1 | B(7)/V(1) |
133 | V* LN Hya | 3497154104039422848 | 194.1256 | -26.4603 | 6.62 | 0.14 | F3Ia | 1/2 | V(2) |
134 | V* LX And | 332909001084177920 | 34.9337 | 40.4562 | 15.71 | 0.07 | - | 1 | V(3) |
135 | V* PS Gem | 3159640386918214528 | 105.9152 | 10.7703 | 7.24 | 0.08 | A0 | 1/2 | B(1,2,7) |
136 | V* RV Col | 2902505745786910080 | 83.9342 | -30.8265 | 8.41 | 0.15 | G5 | 1/2 | V(4) |
137 | V* RX Cap | 6879691160336671744 | 303.7301 | -12.9429 | 11.37 | 0.28 | G0Iae | 1 | V(4) |
138 | V* TT Oph | 4386330497453246080 | 252.3995 | 3.6317 | 9.85 | 0.23 | F5pe | 1 | V(4) |
139 | V* TX Oph | 4392705672029913600 | 256.0004 | 4.9836 | 10.05 | 0.47 | F8Iae | 1 | - |
140 | V* V1027 Cyg | 2030200671149815424 | 300.6141 | 30.0737 | 7.69 | 2.96 | G7Ia | 1/2 | - |
141 | V* V1333 Sco | 6023926760641310208 | 246.5849 | -34.2869 | 10.95 | 1.79 | F8 | 1 | B(7) |
142 | V* V1401 Aql | 4190636669164572928 | 301.2726 | -11.5994 | 6.21 | 0.33 | F2II | 1/2 | - |
143 | V* V2053 Oph | 4470790101628029440 | 273.7058 | 5.2155 | 9.65 | 0.63 | C | 1 | B(1,7) |
144 | V* V340 Ser | 4162959693758887424 | 262.6955 | -11.3689 | 9.3 | 1.31 | F2/3II | 2 | B(7) |
145 | V* V360 Cyg | 1852749557493420032 | 317.6479 | 30.6724 | 11.08 | 0.4 | F8Ie | 1 | - |
146 | V* V399 Cyg | 1869102495165230592 | 312.2852 | 33.6958 | 10.96 | 0.6 | G8 | 1 | V(4) |
147 | V* V400 Sco | 4040579578519065728 | 267.8291 | -36.1824 | 12.7 | 1.7 | - | 1 | - |
148 | V* V421 CMa | 5617989266685365120 | 109.0345 | -23.4504 | 10.49 | 0.95 | F5 | 1/2 | B(1,2,7)/V(1) |
149 | V* V590 Aql | 4222177328438155776 | 304.2856 | -4.0519 | 11.84 | 0.47 | - | 1 | V(4) |
150 | V* V652 Her | 4449366151908979072 | 252.0196 | 13.2618 | 10.51 | 0.1 | - | 1 | R(1) |
151 | V* V709 Car | 5258718997589107712 | 154.8203 | -57.3239 | 9.3 | 1.41 | G8Ia-0 | 1 | B(1,7)/V(4) |
152 | V* V760 Sgr | 4068898810519269248 | 267.5448 | -22.848 | 10.03 | 2.04 | G5 | 1 | - |
153 | V* V802 Car | 5241806275407841664 | 165.518 | -62.1619 | 8.62 | 1.75 | F2III | 1/2 | B(1,2,7)/V(4) |
154 | V* V811 Ara | 5914387846002855296 | 255.9235 | -61.5047 | 10.68 | 0.38 | - | 1 | - |
155 | V* V825 Ara | 5921745812182394496 | 265.0453 | -53.7927 | 11.02 | 0.49 | - | 1 | - |
156 | V* V956 Cen | 6066902993687172608 | 198.5344 | -54.6929 | 7.99 | 0.95 | F5Ia/ab | 1/2 | B(7)/V(4) |
157 | V* YY Ara | 5830750401670303872 | 250.3352 | -59.8752 | 9.47 | 0.58 | K0:-Me | 1 | - |
Num | Distance | Low Dist. | High Dist. | Flag (T) | SED | Flag | |||||||
(pc) | (pc) | (pc) | (K) | () | (kyr) | ||||||||
2 | 1307 | 1199 | 1436 | 0.50 | 6500 | T3(1) | 4.11 | 4.02 | 4.06 | 3.14 | disc | - | |
7 | 1909 | 1870 | 1947 | 3.85 | 4250 | T3(2) | 4.34 | 4.32 | 2.39 | 2.79 | 1.22 | stellar | H |
14 | 6639 | 5922 | 7516 | 1.75 | 16000 | T3(3) | 3.75 | 3.63 | 3.74 | 1.33 | 3.62 | stellar | - |
15 | 2271 | 2160 | 2400 | 4.75 | 3750 | T1 | 3.61 | 3.56 | 2.32 | 1.12 | stellar | H | |
18 | 3467 | 3159 | 3934 | 0.20 | 20000 | T3(4) | 3.55 | 3.44 | 3.47 | 0.93 | 18.14 | shell | H |
19 | 2839 | 2711 | 2962 | 1.60 | 6500 | T3(5) | 3.62 | 3.58 | 3.41 | 1.12 | 5.61 | shell | - |
21 | 4734 | 4150 | 5414 | 0.30 | 21000 | T3(6) | 3.61 | 3.47 | 3.95 | 0.95 | 16.54 | shell | H |
24 | 9965 | 9131 | 11304 | 0.90 | 22000 | T3(7) | 4.30 | 4.19 | 4.15 | 3.81 | 1.10 | stellar | (*) |
25 | 5862 | 5008 | 7121 | 0.73 | 23000 | T3(8) | 4.40 | 4.21 | 4.02 | 3.00 | 9.15 | stellar | H |
26 | 4982 | 4299 | 6049 | 1.00 | 18000 | T3(9) | 4.24 | 4.06 | 3.88 | 3.63 | 0.99 | stellar | - |
27 | 4718 | 4214 | 5732 | 3.30 | 6000 | T2 | 4.12 | 3.95 | 3.98 | 3.16 | disc | - | |
29 | 3858 | 3644 | 4143 | 2.15 | 20000 | T3(6) | 4.43 | 4.37 | 3.09 | 4.00 | 0.69 | shell | - |
30 | 3421 | 3226 | 3635 | 3.14 | 9250 | T2 | 3.87 | 3.81 | 3.66 | 1.98 | 3.36 | shell | - |
31 | 4340 | 4100 | 4599 | 1.97 | 9000 | T2 | 3.77 | 3.72 | 3.63 | 1.42 | 3.94 | shell | - |
33 | 2822 | 2677 | 2982 | 1.78 | 20000 | T3(6) | 4.02 | 3.97 | 3.45 | 2.84 | 2.14 | shell | - |
34 | 5021 | 4607 | 5683 | 1.00 | 8500 | T2 | 3.76 | 3.65 | 3.72 | 1.38 | 3.82 | shell | - |
35 | 10125 | 9128 | 11379 | 2.25 | 6750 | T2 | 3.67 | 3.56 | 3.27 | 0.97 | 9.06 | stellar | H |
37 | 10312 | 8988 | 12281 | 11750 | T2 | 3.45 | 3.28 | 3.70 | 13.83 | shell | H | ||
39 | 2916 | 2712 | 3196 | 0.80 | 7250 | T3(10) | 3.92 | 3.84 | 4.04 | 2.27 | 1.49 | stellar | (*) |
40 | 2099 | 1991 | 2209 | 2.00 | 7250 | T3(11) | 3.94 | 3.89 | 3.37 | 2.40 | 1.63 | shell | H |
41 | 1104 | 1075 | 1139 | 1.00 | 4250 | T3(12) | 4.16 | 4.13 | 4.08 | 3.22 | disc | - | |
42 | 1788 | 1722 | 1845 | 0.82 | 7250 | T3(13) | 3.95 | 3.92 | 3.96 | 2.42 | 1.66 | disc | - |
43 | 1567 | 1503 | 1632 | 0.25 | 8250 | T3(14) | 3.54 | 3.50 | 3.46 | 0.93 | 9.14 | stellar | H |
45 | 5130 | 4530 | 5923 | 0.70 | 7000 | T3(13) | 4.40 | 4.27 | 4.40 | 3.00 | 9.14 | disc | H |
47 | 1708 | 1647 | 1782 | 1.00 | 8250 | T3(11) | 3.74 | 3.71 | 3.63 | 1.32 | 3.57 | shell | - |
49 | 3101 | 2895 | 3388 | 0.60 | 6750 | T3(11) | 4.51 | 4.44 | 4.62 | 3.00 | 9.14 | stellar | H |
51 | 1921 | 1830 | 2016 | 0.75 | 6000 | T3(11) | 3.88 | 3.83 | 3.71 | 2.07 | 0.28 | shell | H |
53 | 1807 | 1726 | 1891 | 0.40 | 9750 | T3(15) | 3.38 | 3.34 | 3.18 | 9.67 | stellar | - | |
54 | 6999 | 5822 | 8254 | 1.40 | 7000 | T3(16) | 4.56 | 4.37 | 4.51 | 3.00 | 9.14 | disc | H |
58 | 2310 | 2165 | 2481 | 1.80 | 8000 | T3(11) | 3.89 | 3.83 | 3.58 | 2.12 | 1.34 | shell | - |
59 | 1390 | 1344 | 1439 | 0.35 | 10000 | T3(17) | 3.57 | 3.54 | 3.41 | 1.07 | 5.54 | stellar | H (*) |
60 | 1410 | 1356 | 1465 | 2.80 | 5250 | T3(11) | 3.94 | 3.90 | 3.44 | 2.35 | 0.58 | shell | - |
62 | 5416 | 5040 | 5792 | 2.19 | 14000 | T3(18) | 3.51 | 3.44 | 3.50 | 10.68 | shell | - | |
63 | 3685 | 3461 | 3942 | 1.80 | 22000 | T3(19) | 4.15 | 4.09 | 2.88 | 3.32 | 0.79 | shell | - |
64 | 4817 | 4587 | 5042 | 3.00 | 5500 | T3(11) | 3.47 | 3.43 | 3.47 | 8.54 | shell | - | |
72 | 5022 | 4555 | 5658 | 3.35 | 9000 | T3(18) | 3.99 | 3.88 | 3.37 | 2.62 | 1.88 | shell | - |
75 | 9082 | 8261 | 10230 | 2.70 | 7500 | T3(11) | 3.88 | 3.78 | 3.61 | 2.08 | 1.29 | shell | - |
98 | 6605 | 5703 | 7346 | 2.25 | 6750 | T2 | 3.38 | 3.25 | 3.12 | 9.14 | disc | - | |
99 | 2082 | 2026 | 2148 | 3.75 | 14000 | T3(20) | 3.91 | 3.88 | 3.31 | 2.23 | 1.45 | shell | - |
102 | 6830 | 5887 | 8163 | 6.75 | 9750 | T3(18) | 3.69 | 3.51 | 3.10 | 1.20 | 8.33 | shell | - |
111 | 10863 | 9455 | 12375 | 1.40 | 15000 | T3(21) | 4.23 | 4.09 | 3.70 | 2.58 | 1.83 | stellar | H |
112 | 7934 | 6777 | 9233 | 1.75 | 19000 | T3(6) | 4.41 | 4.25 | 3.95 | 3.00 | 1.84 | shell | H |
113 | 6437 | 5646 | 7505 | 1.60 | 16000 | T3(22) | 3.44 | 3.29 | 3.20 | 11.32 | shell | H | |
114 | 5354 | 4801 | 6044 | 2.10 | 19000 | T3(23) | 4.01 | 3.89 | 3.40 | 2.77 | 2.07 | shell | - |
116 | 7318 | 6417 | 8628 | 1.05 | 22000 | T3(6) | 4.12 | 3.96 | 3.27 | 2.26 | 1.55 | shell | H |
117 | 5974 | 5146 | 7444 | 1.00 | 5750 | T3(24) | 3.22 | 3.00 | 3.38 | 0.90 | 7.96 | stellar | H |
121 | 7513 | 6290 | 9363 | 0.30 | 17000 | T3(25) | 3.34 | 3.11 | 3.48 | 19.81 | stellar | H | |
127 | 6147 | 5365 | 6967 | 1.20 | 6250 | T3(13) | 3.46 | 3.32 | 3.34 | 8.96 | disc | - | |
128 | 2317 | 2237 | 2410 | 3.50 | 5750 | T3(26) | 3.73 | 3.69 | 3.34 | 1.24 | 8.85 | disc | - |
129 | 4556 | 4246 | 4929 | 1.35 | 6750 | T3(27) | 3.34 | 3.26 | 3.12 | 9.14 | disc | - | |
132 | 10720 | 9413 | 12080 | 1.75 | 8500 | T2 | 4.48 | 4.35 | 4.07 | 3.00 | 9.15 | disc | H |
133 | 1684 | 1619 | 1769 | 1.00 | 6250 | T3(11) | 4.02 | 3.97 | 3.92 | 2.71 | 1.92 | disc | - |
135 | 1813 | 1707 | 1923 | 2.40 | 6000 | T3(28) | 4.49 | 4.43 | 3.63 | 4.00 | 4.05 | disc | - |
136 | 1850 | 1806 | 1893 | 2.25 | 5500 | T2 | 3.80 | 3.78 | 3.83 | 1.62 | 1.53 | shell | - |
137 | 7253 | 6109 | 8872 | 0.70 | 6000 | T2 | 3.40 | 3.17 | 3.46 | 8.29 | stellar | H | |
138 | 2369 | 2250 | 2507 | 2.25 | 6500 | T2 | 3.46 | 3.40 | 2.94 | 8.96 | stellar | - | |
139 | 4682 | 4285 | 5128 | 1.75 | 6250 | T2 | 3.86 | 3.77 | 3.54 | 1.18 | 7.64 | shell | H |
141 | 3948 | 3726 | 4264 | 2.00 | 6250 | T3(29) | 3.37 | 3.31 | 3.35 | 8.96 | disc | - | |
142 | 727 | 714 | 743 | 1.00 | 6750 | T3(11) | 3.47 | 3.45 | 3.25 | 9.14 | shell | - | |
143 | 4188 | 3841 | 4631 | 1.20 | 5000 | T3(27) | 3.67 | 3.58 | 3.65 | 1.18 | 7.13 | disc | - |
144 | 4269 | 3980 | 4580 | 2.00 | 7250 | T2 | 4.10 | 4.03 | 3.91 | 3.13 | 0.67 | disc | - |
145 | 4795 | 4442 | 5245 | 1.50 | 6250 | T2 | 3.41 | 3.32 | 3.11 | 8.96 | stellar | - | |
146 | 4842 | 4502 | 5201 | 2.25 | 5500 | T2 | 3.58 | 3.51 | 3.28 | 1.08 | 4.78 | stellar | - |
147 | 8617 | 7584 | 9718 | 1.70 | 4500 | T2 | 3.34 | 3.21 | 3.40 | 6.83 | stellar | - | |
148 | 5122 | 4745 | 5526 | 10.25 | 7000 | T3(30) | 4.13 | 4.06 | 3.48 | 3.22 | disc | - | |
152 | 2116 | 1995 | 2245 | 2.80 | 5500 | T3(31) | 3.40 | 3.34 | 3.27 | 8.54 | stellar | - | |
153 | 4404 | 4113 | 4693 | 2.00 | 7500 | T3(27) | 4.42 | 4.35 | 4.36 | 4.00 | 9.14 | disc | - |
156 | 1535 | 1468 | 1606 | 1.80 | 6750 | T2 | 3.64 | 3.60 | 3.38 | 1.14 | 6.95 | shell | - |
157 | 2348 | 2251 | 2450 | 2.80 | 4250 | T3(32) | 3.41 | 3.37 | 3.27 | 5.58 | stellar | - |
Temperatures are: (T1) obtained from SED fitting with from Stassun (2019), (T2): derived from Simbad spectral types, (T3): obtained from the literature (mainly for spectroscopic measurements): (1): Luck (2014), (2): Steinmetz et al. (2020), (3): Jeffery et al. (1998), (4): Napiwotzki et al. (1994), (5): Ting et al. (2019), (6): Mello et al. (2012), (7): Venn et al. (1998), (8): Herrero et al. (2020), (9): Jeffery (1993), (10): Rao et al. (2012), (11): Kamath et al. (2022), (12): de Ruyter et al. (2005), (13): Corporaal et al. (2023), (14): Arentsen et al. (2019), (15): Klochkova et al. (2018), (16): Reyniers & Van Winckel (2001), (17): Kodaira & Philip (1984), (18): Raman et al. (2017), (19): Klochkova (2014), (20): Parthasarathy et al. (2000a), (21): Şahin (2018), (22): Parthasarathy et al. (2000b), (23): Ikonnikova et al. (2020), (24): Wallerstein (1958), (25): Drilling et al. (2013), (26): Kluska et al. (2022), (27): Gielen et al. (2011), (28): Waelkens et al. (1991), (29): Maas et al. (2005), (30): Van Winckel (1997), (31): Buder et al. (2021), (32): Henize (1976).
Num | Distance | Low Dist. | High Dist. | T(eff) | Flag (T) | SED | Flag | ||||
(pc) | (pc) | (pc) | (K) | ||||||||
6 | 4007 | 3509 | 4603 | 0.09 | 3500 | T1 | 3.19 | 3.05 | 3.33 | stellar | H |
8 | 7877 | 7285 | 8653 | 7250 | T2 | 2.92 | 2.83 | 3.01 | stellar | H | |
16 | 7633 | 6920 | 8251 | 1.40 | 11000 | T3(1) | 3.12 | 3.03 | 3.20 | stellar | H |
23 | 2569 | 2382 | 2815 | 0.20 | 11750 | T2 | 2.46 | 2.38 | 2.54 | stellar | H |
32 | 2138 | 2058 | 2229 | 0.70 | 9500 | T3(2) | 3.16 | 3.12 | 3.19 | stellar | H |
44 | 2568 | 2429 | 2706 | 0.30 | 7500 | T3(3) | 2.99 | 2.94 | 3.04 | stellar | H |
46 | 4893 | 4514 | 5338 | 0.60 | 6750 | T3(4) | 3.19 | 3.11 | 3.27 | stellar | H |
48 | 1446 | 1391 | 1510 | 0.85 | 22000 | T3(5) | 2.56 | 2.52 | 2.60 | stellar | - |
57 | 945 | 921 | 964 | 0.40 | 6500 | T3(6) | 2.84 | 2.82 | 2.86 | disc | - |
61 | 937 | 926 | 947 | 1.00 | 5750 | T3(7) | 2.06 | 2.05 | 2.07 | shell | - |
66 | 1982 | 1915 | 2056 | 0.66 | 3750 | T3(8) | 2.37 | 2.33 | 2.40 | shell | - |
69 | 1952 | 1895 | 2002 | 0.69 | 3750 | T1 | 2.63 | 2.60 | 2.65 | shell | - |
73 | 5266 | 4836 | 5721 | 4.00 | 1000 | T3(8) | 3.07 | 2.99 | 3.15 | shell | - |
80 | 3226 | 3025 | 3469 | 5.55 | 6000 | T1 | 2.76 | 2.70 | 2.83 | shell | - |
81 | 2504 | 2408 | 2629 | 3.28 | 6750 | T1 | 2.45 | 2.41 | 2.49 | shell | - |
83 | 1978 | 1856 | 2105 | 2.50 | 4000 | T3(9) | 3.01 | 2.95 | 3.06 | disc | - |
86 | 4277 | 3572 | 5089 | 5.80 | 8750 | T3(8) | 3.00 | 2.81 | 3.19 | shell | - |
87 | 5204 | 4054 | 6718 | 2.00 | 3500 | T3(8) | 1.96 | 1.64 | 2.27 | shell | - |
88 | 2533 | 2306 | 2812 | 4.53 | 3750 | T1 | 2.94 | 2.84 | 3.04 | shell | - |
90 | 4882 | 4170 | 5705 | 3.00 | 4250 | T3(8) | 2.27 | 2.10 | 2.43 | shell | - |
91 | 5991 | 5520 | 6694 | 1.76 | 8250 | T1 | 3.01 | 2.91 | 3.11 | shell | - |
100 | 4898 | 4294 | 5685 | 4.79 | 5750 | T1 | 3.01 | 2.87 | 3.16 | disc | - |
101 | 6087 | 5536 | 6736 | 0.67 | 3500 | T3(8) | 3.15 | 3.06 | 3.25 | shell | - |
105 | 1694 | 1631 | 1768 | 5.34 | 4250 | T1 | 2.50 | 2.46 | 2.53 | disc | - |
110 | 5710 | 4998 | 6635 | 14000 | T3(10) | 2.79 | 2.64 | 2.94 | stellar | H | |
130 | 1365 | 1313 | 1420 | 0.90 | 4750 | T2 | 3.21 | 3.16 | 3.27 | shell | - |
149 | 6299 | 5704 | 7065 | 0.47 | 4000 | T3(9) | 3.06 | 2.95 | 3.16 | stellar | H |
154 | 4714 | 4378 | 5080 | 0.38 | 4000 | T1 | 3.29 | 3.22 | 3.36 | stellar | - |
155 | 2123 | 1943 | 2310 | 0.49 | 3500 | T1 | 2.88 | 2.80 | 2.96 | stellar | - |
Temperatures are: (T1) obtained from SED fitting with from Stassun (2019), (T2): derived from Simbad spectral types, (T3): obtained from the literature (mainly for spectroscopic measurements): (1): Mooney et al. (2004), (2): Venn et al. (1998), (3): Kamath et al. (2022), (4): Gonzalez & Wallerstein (1992), (5): Hunger & Kaufmann (1973), (6): Kluska et al. (2022), (7): Ting et al. (2019), (8): Suárez et al. (2006), (9): Steinmetz et al. (2020), (10): Quin & Lamers (1992).
Num | Distance | Low Dist. | High Dist. | T(eff) | Flag (T) | SED | Flag | ||||
(pc) | (pc) | (pc) | (K) | ||||||||
1 | 1139 | 1109 | 1172 | 1.45 | 9250 | T3(1) | 4.28 | 4.25 | 4.30 | stellar | MC |
11 | 3379 | 3021 | 3936 | 0.79 | 5000 | T3(2) | 0.69 | 0.55 | 0.82 | stellar | - |
17 | 307 | 305 | 308 | 0.07 | 6250 | T2 | 1.10 | 1.09 | 1.10 | stellar | - |
38 | 623 | 619 | 628 | 2.75 | 7500 | T3(3) | 1.97 | 1.96 | 1.98 | disc | MC |
50 | 219 | 217 | 220 | 0.50 | 8750 | T2 | 1.40 | 1.39 | 1.40 | stellar | - |
65 | 390 | 387 | 393 | 0.93 | 9500 | T2 | 1.22 | 1.21 | 1.22 | uncertain | - |
84 | 2915 | 2503 | 3473 | 2.20 | 5000 | T3(3) | 0.87 | 0.70 | 1.05 | shell | - |
103 | 1787 | 1746 | 1825 | 2.25 | 5500 | T3(4) | 2.88 | 2.85 | 2.90 | disc | MC |
106 | 1060 | 1046 | 1072 | 1.32 | 6750 | T3(3) | 0.97 | 0.95 | 0.98 | shell | - |
107 | 482 | 478 | 485 | 0.63 | 8500 | T3(2) | 1.38 | 1.37 | 1.39 | disc | - |
108 | 964 | 948 | 981 | 2.05 | 6750 | T3(3) | 0.90 | 0.89 | 0.92 | disc | MC |
109 | 2186 | 1944 | 2596 | 2.53 | 6000 | T1 | 0.72 | 0.56 | 0.87 | shell | - |
134 | 508 | 497 | 520 | 0.07 | 5500 | T1 | -0.94 | -1.01 | -0.87 | disc | - |
151 | 4006 | 3630 | 4403 | 6.75 | 5000 | T2 | 5.32 | 5.23 | 5.42 | disc | MC |
Temperatures are: (T1) obtained from SED fitting with from Stassun (2019), (T2): derived from Simbad spectral types, (T3): obtained from the literature (mainly for spectroscopic measurements): (1): Firnstein & Przybilla (2012), (2): Suárez et al. (2006), (3): Raman et al. (2017), (4): Kluska et al. (2022).
Num | Distance | Low Dist. | High Dist. | T(eff) | Flag (T) | SED | Flag | ||||
(pc) | (pc) | (pc) | (K) | ||||||||
4 | 3506 | 2918 | 4516 | 1.64 | 7250 | T2 | 1.06 | 0.79 | 1.32 | stellar | HB/H |
5 | 4503 | 3936 | 5076 | 0.41 | 9750 | T1 | 1.41 | 1.28 | 1.54 | stellar | HB/H |
20 | 3661 | 3424 | 3917 | 3.75 | 16000 | T2 | 4.69 | 4.62 | 4.75 | disc | S |
22 | 3604 | 3342 | 3931 | 0.06 | 10750 | T1 | 1.78 | 1.7 | 1.86 | stellar | HB/H |
56 | 4432 | 4078 | 4782 | 4.00 | 7500 | T3(1) | 5.47 | 5.4 | 5.55 | shell | S |
140 | 3723 | 3486 | 3988 | 3.8 | 5000 | T2 | 5.13 | 5.06 | 5.19 | disc | S |